# Shaping Sparse Rewards in Reinforcement Learning: A Semi-supervised Approach

### v1

@@ Abstract
In many real-world scenarios, reward signal for agents are exceedingly sparse, making it challenging to learn an effective reward function for reward shaping. To address this issue, our approach performs reward shaping not only by utilizing non-zero-reward transitions but also by employing the Semi-Supervised Learning (SSL) technique combined with a novel data augmentation to learn trajectory space representations from the majority of transitions, zero-reward transitions, thereby improving the efficacy of reward shaping. Experimental results in Atari and robotic manipulation demonstrate that our method effectively generalizes reward shaping to sparse reward scenarios, achieving up to four times better performance in reaching higher best scores compared to curiosity-driven methods. The proposed double entropy data augmentation enhances performance, showcasing a 15.8% increase in best score over other augmentation methods.
1 Introduction
@@
The sparse reward problem is a core challenge in Reinforcement Learning (RL) when solving real-world tasks Kober u. a. (2013). In supervised learning, supervision signals are provided by the training data. In reinforcement learning, rewards take on the role of supervision signals, guiding the agent to optimize its policy. Many real-world tasks naturally have the feature of delayed or infrequent rewards due to the complexity and nature of the tasks. In tasks like Go Silver u. a. (2016), navigation Wang u. a. (2020), or robotic arm manipulation Gu u. a. (2017); Riedmiller u. a. (2018), agents only receive rewards upon successfully achieving the final goal, such as winning a game, reaching a target, or completing a grasp, with no feedback for intermediate steps. Sparse reward provides little immediate feedback to guide the agent’s exploration and high variance in returns, makes it difficult to learn the optimal policy Plappert u. a. (2018b).
@@
A potential solution to tackle the sparse reward problem is reward design and learning Ng und Russell (2000). Methods in this category often operate on the replay buffer, processing off-policy data to implement their respective reward design Schaul u. a. (2015); Andrychowicz u. a. (2017); Peng u. a. (2019). Pathak u. a. (2017) propose the Intrinsic Curiosity Module (ICM) that formulates curiosity as the error in an agent’s ability to predict the consequence of its own actions learned by a self-supervised inverse dynamics model. However, curiosity-driven methods face a significant challenge: curiosity, as an intrinsic reward, is inherently decoupled from the actual objectives and tasks. This may cause agents to overly focus on “new” but meaningless states. Kumar u. a. (2019) propose Reward Conditioned Policy (RCP) that treats non-expert trajectories collected from sub-optimal policies as optimal supervision for matching the reward of a given trajectory. However, methods based on suboptimal policies under supervised learning suffer from poor sample efficiency Peng u. a. (2019), especially when non-zero-reward transitions are extremely rare in sparse reward cases.
Moreover, there exists alternative approaches beyond self-supervised methods that may better address the challenges posed by sparse reward data. Raileanu u. a. (2020) utilize data augmentation to enhance RL algorithm, yet such augmentation methods are restricted to image data and tend to lose or distort information in non-vision-based task, as the nature of image data is inherently different from that of time-series data, text data, or feature vector data.
In this paper, our approach performs reward shaping not only by utilizing non-zero-reward transitions but also by employing Semi-Supervised Learning (SSL) technique to learn trajectory space representations from the majority of transitions, i.e., zero-reward transitions, thereby improving reward shaping, it could help bridge the gap in handling sparse rewards.
@@
Explicitly, we propose an RL algorithm that applies the idea of SSL, called Semi-Supervised Reward Shaping framework (SSRS), and also propose a new data augmentation technique double entropy data augmentation specifically for non-image-based task as part of the RL pipeline. By progressively approximating the state-value function and the action-state function, we shape the reward in sparse-reward trajectories. Consistency regularization to both zero-reward and non-zero-reward transitions is adopted for the optimization of the reward estimator. Additionally, a monotonicity constraint is incorporated over the two components of the reward shaping estimator to further reduce the discrepancy between the shaped reward distribution and the true reward distribution.
@@
We evaluate the performance of SSRS in reward-sparse Atari and robotic manipulation environments, comparing it with ICM Pathak u. a. (2017), RCP Kumar u. a. (2019), Prioritized Experience Replay (PER , Schaul u. a. (2015)). Our model demonstrates performance comparable to ICM in Atari environments and achieves a maximum increase over ICM by a factor of four in reaching higher best scores. We further validate that the proposed double entropy data augmentation enhances performance showcasing a 15.8% increase in best score compared to other augmentation methods.
2 Related Work
Currently, research aiming at addressing the sparse reward problem mainly focuses on the following areas: experience replay mechanisms Peng u. a. (2019) and reward design and learning. Schaul u. a. (2015) develop a framework for prioritizing experience, so as to replay important transitions more frequently, and therefore learn more efficiently. Memarian u. a. (2021) leverage self-supervised methods to extract signals from trajectories while simultaneously updating the policy. Andrychowicz u. a. (2017) propose Hindsight Experience Replay (HER) that allows sample-efficient learning from rewards which are sparse and binary and therefore performs well especially on robotic manipulation tasks. Ng und Russell (2000) propose the idea of learning reward functions from optimal interaction sequences, known as Inverse Reinforcement Learning (IRL). Brown u. a. (2019) propose an IRL algorithm that outperforms the demonstrator by extrapolating beyond a set of ranked demonstrations in order to infer high-quality reward functions.
@@ 2 Related Work
Besides expert trajectories, reward shaping can also be applied based on different evaluation criteria. Count-based methods Choi u. a. (2018); Ostrovski u. a. (2017); Bellemare u. a. (2016) incentivize agents based on the rarity of states, while curiosity-driven exploration methods reward the agent’s exploratory behavior. ICM Pathak u. a. (2017) defines curiosity as the error in the agent’s ability to predict the consequences of its own actions in a visual feature space. This curiosity signal serves as an intrinsic reward, driving the agent to explore.
Kumar u. a. (2019) employ supervised learning techniques, viewing non-expert trajectories collected from sub-optimal policies as optimal supervision for matching the reward of the given trajectory. Peng u. a. (2019) simplify the process into regressing target values for a value function, and weighted target actions for the policy.
@@
Another related direction is the application of data augmentation Shorten und Khoshgoftaar (2019) in RL algorithms Kostrikov u. a. (2020). Hansen und Wang (2020) decouple augmentation from policy learning and is found to significantly advance sample efficiency, generalization, and stability in training over vision-based RL methods. Yarats u. a. (2021) use data augmentation to learn directly from pixels and is able to solve complex humanoid locomotion tasks directly from pixel observations. Studies like Lin u. a. (2019) generate feasible trajectories based on symmetries observed in the trajectory space for robot control tasks, thereby constructing an Invariant Transform Experience Replay framework to address the issue of high sample requirements. Raileanu u. a. (2020) explore the use of data augmentation techniques in visual RL tasks, such as random cropping, random noise, and color jittering, to improve the robustness of the policy $\pi(a|s)$ and the value function $V(s)$.
3 Methodology
In this section, we introduce the proposed method by first introducing the proposed double entropy data augmentation, followed by the SSRS framework, featuring a monotonicity constraint and consistency regularization.
3.1 Double Entropy Data Augmentation
In reinforcement learning, applying invariant transformations to trajectories can improve the generalization performance of both the policy network and the value function network Lin u. a. (2019); Raileanu u. a. (2020). The unitless property of Entropy makes it particularly useful in data augmentation, as it allows for a fair comparison of information content across different features, regardless of their scale or domain. Entropy is essentially a measure of uncertainty or randomness in a probability distribution, and it is calculated in a way that normalizes the result, making it dimensionless. We propose a double entropy data augmentation method as follow. Shannon Entropy Shannon (1948) quantifies the amount of information required to describe or encode a random variable’s possible states. In RL tasks, observations and trajectories are matrices in practice, but can also be normalized and viewed as discrete random variable. Let $X$ be a discrete random variable with finite sample space $\mathcal{X}$ and probability mass function $p(x)=$Pr$\left\{X=x\right\}$, $x\in\mathcal{X}$. The entropy $H(X)$ of a discrete random variable $X$ is defined by
$H(X):=-\sum_{x\in\mathcal{X}}p(x)\log p(x).$
Consider a $m\times n$ matrix $A$ with each element $a_{ij}\leq 0,\,\forall i=1,...,m,\,j=1,...,n$, and the Shannon Entropy of matrix $A$ is defined by
$H(A):=-\sum_{k=1}^{mn}p_{k}\log(p_{k}),$ (1)
Denote state space as $\mathcal{S}\in\mathbb{R}^{m_{1}}$ and action space as $\mathcal{A}\in\mathbb{R}^{m_{2}}$, with $m_{1},\,m_{2}\geq 1$. Consider a trajectory $\tau$ of length $N$ as $\langle s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},\dots,s_{N}\rangle$ with $a_{t}\in\mathcal{A},\ s_{t}\in\mathcal{S}$, for all $t=1,\dots,N$, and rewards are given following the environment’s original reward function $r_{t}=R(s_{t},a_{t},s^{\prime}_{t})$. The set of all trajectories can be stacked into the matrix $\bar{\Gamma}=\begin{bmatrix}S,\,A,\,R\end{bmatrix}$ with $S\in\mathbb{R}^{N\times m_{1}},\,A\in\mathbb{R}^{N\times m_{2}}$ and $R\in\mathbb{R}^{N\times 1}$, and now we define the transform double entropy data augmentation $\sigma$ over matrix $\bar{\Gamma}$ as:
where $h^{n}\in\mathbb{R}$, equals to the entropy $H(s^{n})$, matrix $s^{n}$ is a submatrix of the stacked state $S$, obtained by equally dividing $S$ into $n$ parts along the state dimension. Here $n$ is a hyperparameter and $H(s)$ is the matrix Shannon entropy defined in Eq.(1).
3.2 Semi-Supervised Reward Shaping
@@
The Semi-Supervised Reward Shaping (SSRS) framework proposed in this paper, aims to fit an optimal state-value function $V^{*}(s)$ and an optimal action-value function $Q^{*}(s,a)$, which are used to shape the rewards of trajectory $\Gamma$. Given an one-step trajectory $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$, the estimated reward can be calculated from the estimations of the optimal state-value function $V^{*}(s)$ and the optimal action-value function $Q^{*}(s,a)$ as follows. We first define the confidence score vector of timestep $t$ over a fixed reward set $Z=\{z_{i},\ i=1,\dots,N_{z}\}$ in Eq.(2), where $N_{z}$ is a hyperparameter controlling the number of estimated rewards. Value of $z_{i}$ is set according to the collected true reward value, and is updated throughout the process of agent’s interactions with environment.
@@
By the following Eq.(2), we can get confidence vector $\mathbf{q_{t}}\in\mathbb{R}^{N_{z}}$ at timestep $t$, with each $q_{i}\in\mathbf{q_{t}}$ corresponding to a estimated reward $z_{i}\in Z$, computed as
@@
A Monotonicity Constraint in SSRS
The value function $V(s_{t})$ provides a global baseline for a state, while advantage function captures the relative advantage or disadvantage of action $a$ with respect to this baseline. This separation makes it easier to assess the relative importance of specific actions, and thus compensates for the insensitivity to actions because the disturbance of data augmentation only takse place on state $s$. By utilizing the relationship between $V$ and $Q$, we can introduce the monotonicity constraint in the SSRS framework as follows, which quantifies the mean square positive advantage function values. Define $\delta_{t}(\theta)=Q(s_{t},a_{t},\theta_{1})-V(s_{t},\theta_{2})$ where $\theta$ is the combination of $\theta_{1}$ and $\theta_{2}$, we have
@@ A Monotonicity Constraint in SSRS
Note that $\mu$ represents the proportion of non-zero rewards in the buffer, which quantifies the sparsity level, and $B$ is the batch size of trajectories. Minimizing the objective loss $L_{QV}$ over the parameter $\theta$ of can achieve a more stable update, which will be validated in Section 4.3, and is crucial to SSRS framework’s success.
B Consistency Regularization in SSRS
@@
Consistency regularization is an important component of algorithms in the field of SSL Bachman u. a. (2014). The reward estimator, as a crucial component in reward shaping, needs to capture the invariance of trajectories, which refers to the reward unaffected by minor variations in input transitions, aligning with the characteristics of consistency regularization. Furthermore, in scenarios with sparse rewards, applying weak strong augmentations to transitions with zero and non-zero reward transitions before conducting consistency regularization further enhances the generalization ability of the reward estimator in sparse reward environments. Note that the vector $\mathbf{q}_{t}$, the confidence of the reward predictions under the trajectory $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$, consists of $Q(s_{t},a_{t},\,\theta_{1})$ and $V(s_{t+1},\,\theta_{2})$. To be explicit, we use $\mathbf{q}_{t}(\theta)$ to show its dependence in $\theta$. In order to obtain the optimal $\theta$ value for accurately estimating the reward value of each trajectory $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$, SSRS framework first optimizes the following objective loss function:
$L_{r}=\frac{1}{\mu B}\sum_{t=1}^{\mu B}\mathbbm{1}(\max(\mathbf{q}_{t}(\theta))\geq\lambda)\left({r_{t}}-\mathbf{q}_{t}(\theta)\right)^{2}.$ (4)
@@
SSRS also computes the loss on sparse reward trajectories as shown in Eq.(5), which is the loss between the strong augmentation term and weak augmentation term. Assuming continuity of trajectories in the metric space, it calculates the confidence of reward values after weak and strong augmentations, denoted by $\mathbf{q_{t}^{w}}$ and $\mathbf{q_{t}^{s}}$, respectively. Here we omit their dependence on $\theta$ for simplicity. The weak and strong augmentations refer to Table 5 in Appendix, where we consider Gaussian noise with smaller parameters as weak augmentation, and other augmentation such as smoothing, translation, and cutout, with larger parameters as strong augmentation. For confidence values greater than the threshold in weak augmentation, the one-hot operation is performed Bachman u. a. (2014), denoted as $\mathbf{q_{t}^{w,\,\prime}}$, and then used to compute the cross-entropy loss with the normalized confidence values greater than the threshold in strong augmentation, denoted as $\mathbf{q_{t}^{s,\,\prime}}$. The loss function is denoted as.
$L_{s}=\frac{1}{(1-\mu)B}\sum_{t=1}^{(1-\mu)B}$ $[\mathbbm{1}(\max(\mathbf{q_{t}^{s}})\geq\lambda,\,\max(\mathbf{q_{t}^{w}})\geq\lambda)$
@@
where $\mathcal{H}$ denotes the cross-entropy, $-\langle\mathbf{q_{t}^{w,\,\prime}},\,\log(\mathbf{q_{t}^{s,\,\prime}})\rangle$. Together, $L_{r}$ and $L_{s}$ forms the consistency regularization in the SSRS framework.
C Overall Framework
@@
The illustration of the SSRS framework is shown in Figure 1. We also present the value-based SSRS framework with synchronous update, i.e., value function update and reward shaping carry out simultaneously. in Algorithm 1. The loss function of the Semi-Supervised Reward Shaping framework is given by Eq.(3), (4) and (5) above, and combined by the parameter $\alpha$. After the agent collects experience, the loss function can be back-propagated either synchronously or asynchronously to initialize the reward estimation network, and the weight of $L_{s}$ is initialized with a small $\alpha$. It is important to note that the reward estimation network consists of both the $V$ network and the $Q$ network, and during backpropagation, they are treated as a single entity for gradient computation using the chain rule. In Appendix, we provide the derivations of the “approximate” gradients of loss functions $L_{QV}$, $L_{r}$ and $L_{s}$ with respect to $\theta$, after approximating the indicator functions by sigmoid function.
Algorithm 1 Value-based SSRS Framework (synchronous update)
@@
1:Confidence threshold $\lambda$, loss weight $\alpha$, set $Z$ with size $N_{z}$, episodes $T$, update probability $p_{u}$;
2:Initialize parameters $\theta$, $Q$ function, replay buffer $\mathcal{D}$;
3:for $t=1$ to $T$ do
4:     repeat
@@
5:         Take $a_{t}$ from $s_{t}$ using policy derived from $Q$ (e.g., $\epsilon$-greedy), observe $r_{t}$, $s_{t+1}$
6:         if $r_{t}\notin\{r_{1},\dots,r_{t-1}\}$ then
7:              Obtain $N_{z}$ new $z_{i}$ by interpolating from all historical observation $r_{1},...,r_{t}$
8:         end if
9:         if $r_{t}=0$ then
@@
10:              Calculate confidence $\mathbf{q_{t}}(\theta)$ over set $Z$ under parameter $\theta$.
11:              Perform reward shaping on zero-reward transitions in $\mathcal{D}$ of the $p_{u}$ ratio.
12:         end if
@@
13:         Choose a batch of transitions $\mathcal{B}$ from $\mathcal{D}$. Update $Q$ function with shaped reward $\hat{r}$,

### v2

@@ Abstract
In many real-world scenarios, reward signal for agents are exceedingly sparse, making it challenging to learn an effective reward function for reward shaping. To address this issue, the proposed approach in this paper performs reward shaping not only by utilizing non-zero-reward transitions but also by employing the Semi-Supervised Learning (SSL) technique combined with a novel data augmentation to learn trajectory space representations from the majority of transitions, i.e., zero-reward transitions, thereby improving the efficacy of reward shaping. Experimental results in Atari and robotic manipulation demonstrate that our method outperforms supervised-based approaches in reward inference, leading to higher agent scores. Notably, in more sparse-reward environments, our method achieves up to twice the peak scores compared to supervised baselines. The proposed double entropy data augmentation enhances performance, showcasing a 15.8% increase in best score over other augmentation methods
Introduction
@@
The sparse reward problem is a core challenge in Reinforcement Learning (RL) when solving real-world tasks (Kober, Bagnell, and Peters 2013). In supervised learning, supervision signals are provided by the training data. In reinforcement learning, rewards take on the role of supervision signals, guiding the agent to optimize its policy. Many real-world tasks naturally have the feature of delayed or infrequent rewards due to the complexity and nature of the tasks. In tasks like Go (Silver et al. 2016), navigation (Wang et al. 2020), or robotic arm manipulation (Gu et al. 2017; Riedmiller et al. 2018), agents only receive rewards upon successfully achieving the final goal, such as winning a game, reaching a target, or completing a grasp, with no feedback for intermediate steps. Sparse reward provides little immediate feedback to guide the agent’s exploration and high variance in returns, makes it difficult to learn the optimal policy (Plappert et al. 2018).
@@
A potential solution to tackle the sparse reward problem is reward design and learning (Ng, Harada, and Russell 1999). Methods in this category often operate on the replay buffer, processing off-policy data to implement their respective reward design (Schaul et al. 2015; Andrychowicz et al. 2017; Peng et al. 2019). Pathak et al. (2017) propose the Intrinsic Curiosity Module (ICM) that formulates curiosity as the error in an agent’s ability to predict the consequence of its own actions learned by a self-supervised inverse dynamics model. However, curiosity-driven methods face a significant challenge: curiosity, as an intrinsic reward, is inherently decoupled from the actual objectives and tasks. This may cause agents to overly focus on “new” but meaningless states. Kumar, Peng, and Levine (2019) propose Reward Conditioned Policy (RCP) that uses supervised learning as an incentive, thereby eliminating reliance on manually designed novelty as the objective. Building upon this, Self-Supervised Online Reward Shaping (SORS, Memarian et al. (2021)) formalizes a self-supervised learning framework and apply it to sparse-reward environments. In SORS, the original sparse reward provides a self-supervisory signal for reward inference by ranking trajectories that the agent observes, while the policy update is performed with the newly inferred, typically dense reward function. However, methods based on non-expert trajectories under supervised learning suffer from poor sample efficiency (Peng et al. 2019), especially when non-zero-reward transitions are extremely rare in sparse reward cases.
@@
In order to bridge the gap in handling sparse rewards, our approach performs reward shaping not only by utilizing non-zero-reward transitions but also by employing Semi-Supervised Learning (SSL) technique to learn trajectory space representations from the majority of transitions, i.e., zero-reward transitions, thereby improving reward shaping. Explicitly, we propose an RL algorithm that applies the idea of SSL, called Semi-Supervised Reward Shaping framework (SSRS), and also propose a new data augmentation technique called double entropy data augmentation applied on non-image transitions as the perturbation method for SSL. By progressively approximating the state-value function and the action-state function in the reward estimator, we shape the reward in sparse-reward trajectories. Consistency regularization to both zero-reward and non-zero-reward transitions is adopted for the optimization of the reward estimator. Additionally, a monotonicity constraint is incorporated over the two components of the reward shaping estimator to further reduce the discrepancy between the shaped reward distribution and the true reward distribution.
@@
We evaluate the performance of SSRS in reward-sparse Atari and robotic manipulation environments, comparing it with SORS (Memarian et al. 2021), RCP (Kumar, Peng, and Levine 2019) methods. Our model demonstrates performance superior than other supervised-based approaches and achieves a maximum increase over RCP by a factor of two in reaching higher best scores. We further validate that the proposed double entropy data augmentation enhances performance showcasing a 15.8% increase in best score compared to other augmentation methods.
Related Work
@@
There are different reward shaping methods for different objectives: Count-based methods (Choi et al. 2019; Ostrovski et al. 2017; Bellemare et al. 2016) incentivize agents based on the rarity of states, while curiosity-driven exploration methods reward the agent’s exploratory behavior. ICM Pathak et al. (2017) also pursues rarity, but operationalizes it by defining curiosity as the error in the agent’s ability to predict the consequences of its own actions in a visual feature space. This curiosity signal serves as an intrinsic reward, driving the agent to explore novel states. Andrychowicz et al. (2017) propose Hindsight Experience Replay (HER) that enables learning from failures by treating unachieved goals as alternative goals for experience replay.
Approaches using supervised learning for reward shaping don’t rely on specific objectives. The study (Kumar, Peng, and Levine 2019) employs supervised learning techniques, viewing non-expert trajectories collected from sub-optimal policies as optimal supervision to do the reward shaping. Memarian et al. (2021) leverage self-supervised methods to extract signals from trajectories while simultaneously updating the policy.
We extend supervised learning to a semi-supervised paradigm and introduce SSL techniques into the RL framework, specifically employing consistency regularization. and pseudo-labeling (Bachman, Alsharif, and Precup 2014) in the trajectory space. In the realm of SSL, consistency regularization leverages unlabeled data by capitalizing on the invariance assumption—that the model should yield consistent predictions when presented with perturbed variants of identical inputs. Berthelot et al. (2019); Sohn et al. (2020) further improve this methodology on image classification tasks. Pseudo-labeling (Lee 2013) utilizes the model’s predicted hard labels (obtained via $\arg\max$ operations) for unlabeled data, retaining only those predictions whose confidence scores $\mathbf{q}$ exceed the threshold $\tau$, and optimizes them via:
$\mathcal{L}_{u}=\mathds{1}(\max(\mathbf{q})\geq\tau)\cdot\mathcal{H}(\mathbf{q},\arg\max(\mathbf{q})).$ (1)
Note we directly write $\mathbf{q}$, omitting the dependence of $\mathbf{q}$ on the parameters and $\mathds{1}(\cdot)$ is the indicator function.
@@
Another related direction is the application of data augmentation for trajectories (Shorten and Khoshgoftaar 2019) in RL algorithms (Yarats, Kostrikov, and Fergus 2021). Most of them process image data to enhance the agent’s generalization performance. Hansen and Wang (2020) decouple augmentation from policy learning and is found to significantly advance sample efficiency, generalization, and stability in training over vision-based RL methods. Yarats et al. (2022) use data augmentation to learn directly from pixels and is able to solve complex humanoid locomotion tasks directly from pixel observations. Studies like Lin et al. (2019) generate feasible trajectories based on symmetries observed in the trajectory space for robot control tasks, thereby constructing an Invariant Transform Experience Replay framework to address the issue of high sample requirements.
Methodology
Figure 1: The figure illustrates how the SSRS method leverages both non-zero reward transitions and sparse reward transitions for reward shaping through a semi-supervised learning (SSL) approach incorporating strong-weak data augmentation. The augmentation techniques include Cutout, Smooth, and Double Entropy Augmentation, among others (see Table 2 in the Appendix for details).
In this section, we introduce the proposed method and its pipeline is graphically illustrated in Figure 1, which consists of two components: a reward estimator optimized through semi-supervised learning, and a reward estimator optimized through semi-supervised learning and the data augmentation technique for zero-reward transitions that serves as the perturbation required for semi-supervised learning. We will elaborate the key parts of the method in the following subsections and the detailed steps are shown in Algoirthm 1.
Semi-Supervised Reward Estimator
@@
The reward estimator is the very module that takes (augmented) transitions as input and outputs estimated rewards (see Figure 1). During the training phase, the estimator is optimized via consistency regularization using the cross-entropy. In the prediction phase, the final estimated reward is generated through $\arg\max$ selection, choosing the $z$ value corresponding to the peak of the confidence score distribution. The SSRS framework proposed in this paper, aims to fit an optimal state-value function $V^{*}(s)$ and an optimal action-value function $Q^{*}(s,a)$, which are used to shape the rewards of trajectory. Note we use $Q(s,a)$ and $V(s)$ to denote aforementioned function before parameter optimization. Given an one-step trajectory $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$, the estimated reward can be calculated from the estimations of the state-value function $V(s)$ and the action-value function $Q(s,a)$ as follows. We first define the confidence score vector of timestep $t$ over a fixed reward set $Z=\{z_{i},\ i=1,\dots,N_{z}\}$ in Eq.(2), where $N_{z}$ is a hyperparameter controlling the number of estimated rewards. The value of fixed reward set is initialized according to the collected true reward value, and is updated throughout the process of agent’s interactions with environment.
@@
By the following Eq.(2), we can get confidence vector $\mathbf{q_{t}}\in\mathbb{R}^{N_{z}}$ at timestep $t$, with each element $q_{i}\in\mathbf{q_{t}}$ corresponding to a estimated reward $z_{i}\in Z$, computed as
@@
Note that we will use $\mathbf{q}_{t}(\theta)$ in subsequent sections to explicitly denote its dependence on $\theta$, which consists of $\theta_{1}$ (the parameter of the $Q$ function) and $\theta_{2}$ (the parameter of the $V$ function). During the prediction stage, we apply shaping to a proportion $p_{u}$ of transitions in the buffer, where $p_{u}$ dynamically adjusts throughout training. Specifically, $p_{u}$ is initially set proportional to $\log(\mathcal{N}_{r})$, then increased to match $\mathcal{N}_{r}$ during the mid-training phases, before being reduced again to scale with $\log(\mathcal{N}_{r})$. Here, $\mathcal{N}_{r}$ represents the count of trajectories containing genuine reward signals. This prevents excessive variance in the initial policy gradient updates or $Q$-network updates, while in later stages, phases out reliance on shaped rewards as the policy converges toward optimal behavior.
In order to obtain the optimal $\theta$ value for accurately estimating the reward value of each trajectory $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$, SSRS framework first optimizes the following loss function using only transitions with non-zero rewards, so as to minimize the discrepancy between the reward estimator’s predictions and the ground truth values for these transitions. The handling of zero-reward transitions will be addressed in consistency regularization section.
$L_{r}=\frac{1}{\mu B}\sum_{t=1}^{\mu B}\mathds{1}(\max(\mathbf{q}_{t}(\theta))\geq\lambda)\left({r_{t}}-\alpha(\mathbf{q_{t}}(\theta),\lambda)\right)^{2}.$ (3)
Note that $\mu$ represents the proportion of non-zero rewards in the buffer, which quantifies the sparsity level, and $B$ is the batch size of trajectories.
Monotonicity Constraint
Since the estimator employs two neural networks, we introduce a constraint to regulate their outputs and stabilize the training process. The value function $V(s_{t})$ provides a global baseline for a state, while advantage function captures the relative advantage or disadvantage of action $a$ with respect to this baseline. This separation makes it easier to assess the relative importance of specific actions, and thus compensates for the insensitivity to actions because the disturbance of data augmentation only takse place on state $s$. By utilizing the relationship between $V$ and $Q$, we can introduce the monotonicity constraint in the SSRS framework as follows, which quantifies the mean square positive advantage function values. Define advantage function $\delta_{t}(\theta)=Q(s_{t},a_{t},\theta_{1})-V(s_{t},\theta_{2})$ where $\theta$ is the combination of $\theta_{1}$ and $\theta_{2}$, we have
@@
Minimizing the objective loss $L_{QV}$ over the parameter $\theta$ of can achieve a more stable update, which will be validated in experiment section, and is crucial to SSRS framework’s success.
Consistency Regularization
@@
Consistency regularization is a paradigm in the field of SSL (Sajjadi, Javanmardi, and Tasdizen 2016), and is also the key to optimize the reward estimator with the limited feedback reward signal in the sparse reward trajectories. The reward estimator, as a crucial component in SSRS, needs to capture the invariance of trajectories, which refers to the estimated reward unaffected by minor variations in input transitions, aligning with the characteristics of consistency regularization. Furthermore, in scenarios with sparse rewards, applying weak strong augmentations to transitions with zero and non-zero reward transitions before conducting consistency regularization further enhances the generalization ability of the reward estimator in sparse reward environments.
The weak and strong augmentations refer to Table 2 in Appendix, where we consider Gaussian noise with smaller parameters as weak augmentation, and other augmentation such as smoothing, translation, and cutout, with larger parameters as strong augmentation. In conventional data augmentation for weak-strong augmentation, standard techniques such as cutout, flipping, and rotation are typically employed. The unitless property of Entropy makes it particularly useful in data augmentation, as it allows for a fair comparison of information content across different features, regardless of their scale or domain. Entropy is essentially a measure of uncertainty or randomness in a probability distribution, and it is calculated in a way that normalizes the result, making it dimensionless. Building on this intuition, we further adapt a novel double entropy data augmentation method to serve in SSRS on non-image tasks. The implementation details are provided in Appendix.
@@
More specifically, for transitions with zero rewards, the reward estimator computes the loss as shown in Eq.(5), which is the loss between the strong augmentation term and weak augmentation term following typical loss design in consistency regularization. Assuming continuity of trajectories in the metric space, it calculates the confidence of reward values after weak and strong augmentations, denoted by $\mathbf{q_{t}^{w}}$ and $\mathbf{q_{t}^{s}}$, respectively. The calculation of $\mathbf{q_{t}^{w}}$ and $\mathbf{q_{t}^{s}}$ follows Eq.(2), simply replacing the input state s with its weakly or strongly augmented version. Here we omit their dependence on $\theta$ for simplicity. For confidence values greater than the threshold in weak augmentation, the one-hot operation is performed (Bachman, Alsharif, and Precup 2014), denoted as $\mathbf{q_{t}^{w,\,\prime}}$, and then used to compute the cross-entropy loss with the normalized confidence values greater than the threshold in strong augmentation, denoted as $\mathbf{q_{t}^{s,\,\prime}}$. The loss function is denoted as.
$L_{s}=\frac{1}{(1-\mu)B}\sum_{t=1}^{(1-\mu)B}$ $[\mathds{1}(\max(\mathbf{q_{t}^{s}})\geq\lambda,\,\max(\mathbf{q_{t}^{w}})\geq\lambda)$
@@
where $\mathcal{H}$ denotes the cross-entropy.
@@
The algorithm for value-based SSRS with synchronous update—i.e., value function update and reward shaping carry out simultaneously— is outlined in Algorithm 1. The agent begins by collecting experience transitions. The reward estimation network is then initialized through back-propagation, which can be performed either synchronously or asynchronously using the loss function. The complete SSRS loss function, as specified in Eqs. (3), (4) and (5), combines these components through the weighting parameter $\alpha$. In Appendix, we provide the derivations of the “approximate” gradients of loss functions $L_{QV}$, $L_{r}$ and $L_{s}$ with respect to $\theta$ in Appendix, after approximating the indicator functions by sigmoid function.
Algorithm 1 Value-based SSRS Framework
@@
Parameter Confidence threshold $\lambda$, loss weight $\alpha$, set $Z$ with size $N_{z}$, episodes $T$, update probability $p_{u}$;
1: Initialize parameters $\theta$, $Q$ function, replay buffer $\mathcal{D}$;
2:for $t=1$ to $T$ do
3:  repeat
@@
4:   Take $a_{t}$ from $s_{t}$ using policy derived from $Q$ (e.g., $\epsilon$-greedy), observe $r_{t}$, $s_{t+1}$
5:   if $r_{t}\notin\{r_{1},\dots,r_{t-1}\}$ then
6:    Obtain $N_{z}$ new $z_{i}$ by interpolating from all historical observation $r_{1},...,r_{t}$
7:   end if
8:   if $r_{t}=0$ then
@@
9:    Calculate confidence $\mathbf{q_{t}}(\theta)$, $\mathbf{q_{t}^{w,\,\prime}(\theta)}$ of non-zero and zero reward transitions over set $Z$ under parameter $\theta$.
10:    Perform reward shaping on zero-reward transitions in $\mathcal{D}$ of the $p_{u}$ ratio.
11:   end if
@@
12:   Choose a batch of transitions $\mathcal{B}$ from $\mathcal{D}$ and update $Q$ function with shaped reward $\hat{r}$,